David currently has a square garden.he wants to redesign his garden and make it into a rectangle with a length that is 3 feet shorter than twice its width.he decides that the perimeter should be 60 feet.what is the dimensions,in feet,of his new garden.
Answer by richwmiller(17219)

2L+2W=60
L+W=30,
L=2W-3
2W-3+W=30
3W=33
W=11
L=19
check
22-3=19
ok